# Data Publishing in Mechanics and Dynamics: Challenges, Guidelines, and Examples from Engineering Design

### v1

@@ Abstract
Data-based methods have gained increasing importance in engineering, especially but not only driven by successes with deep artificial neural networks. Success stories are prevalent, e.g., in areas such as data-driven modeling, control and automation, as well as surrogate modeling for accelerated simulation. Beyond engineering, generative and large-language models are increasingly performing and helping with tasks that, previously, were solely associated with creative human processes. Thus, it seems timely to seek artificial-intelligence-support for engineering design tasks to automate, help with, or accelerate purpose-built designs of engineering systems, e.g., in mechanics and dynamics, where design so far requires a lot of specialized knowledge. However, research-wise, compared to established, predominantly first-principles-based methods, the datasets used for training, validation, and test become an almost inherent part of the overall methodology. Thus, data publishing becomes just as important in (data-driven) engineering science as appropriate descriptions of conventional methodology in publications in the past. This article analyzes the value and challenges of data publishing in mechanics and dynamics, in particular regarding engineering design tasks, showing that the latter raise also challenges and considerations not typical in fields where data-driven methods have been booming originally. Possible ways to deal with these challenges are discussed and a set of examples from across different design problems shows how data publishing can be put into practice. The analysis, discussions, and examples are based on the research experience made in a priority program of the German research foundation focusing on research on artificially intelligent design assistants in mechanics and dynamics.
@@ 1 Introduction
Secondly, the article presents a set of six examples from the aforementioned research projects that outline how the described challenges of data publication were overcome. The authors hope that these concrete examples, all based on published, state-of-the-art research, will further nurture ongoing discussions and change within the community and will be of help to other scientists in mechanics and dynamics who want to publish data and data-driven research.
Thirdly, while not the immediate goal, the article and its examples may show to scientists from outside engineering that very challenging and worthwhile problems exist in mechanics, dynamics, and engineering design, where good data-driven contributions could lead to sustained technological progress. The authors invite data scientists to try their methods on the data published in the publications associated with the described examples.
@@
The article is organized as follows. Based on a general motivation in Section 2, challenges for publishing data from the perspective of the priority program are formulated in Section 3. Ideas and guidelines how these challenges might be addressed are discussed in Section 4. Section 5 presents examples of published datasets, program code, and model problems from within the priority program. Finally, Section 6 draws conclusions and gives and outlook with regard to the findings specific for engineering design.
@@ 2 The Value of Publishing Data
It is to be expected that the availability of data and well-prepared benchmark problems can very much influence and steer in which areas progresses in machine learning are made. However, an analysis of published datasets and benchmark problems shows that there are comparatively few from the field of mechanics and dynamics, especially when focusing on published material suitable for people that are not experts in the field.
Obviously, apart from the kind of data mentioned before, there have always been published datasets that are not interesting primarily due to their value as a basis for further processing. For instance, measurements of temperatures at a certain geographic location or measurements of certain material properties are among these datasets for which the provision of the data itself is the purpose of the investigation and publication. This kind of data publishing is not the main purpose of this paper.
@@ 3 Challenges in Data Publishing
Some challenges to be faced for the publication of data are quite common beyond the specific characteristics of the priority program. These include:
  • Interpretability and Domain Knowledge: Raw data can be tough to understand for anybody unfamiliar with the generation process of the data. Furthermore, the representation of the data may heavily depend not only on the underlying system but also on the chosen analysis method and observation approach. In simulations, the deformation of a simple mechanical beam, for example, can be computed using numerical discretization schemes that involve the partitioning of the physical domain into small elements. The shape and/or numbering of those elements may however greatly vary across different simulation approaches, while displaying the same or very similar elastic deflection of the beam. Comparability of different simulation techniques – and the resulting data – can be difficult for the very same physical system. This can happen analogously in experiments, if measurements are taken at different points of the same mechanical object or with different measurement hardware, which may come with different built-in filtering, e.g., for noise reduction. In the light of this challenge, it is maybe not surprising that many of the first, renewed successes of machine learning in recent years happened for data interpretable without special education, like text, speech, and image data.
@@
  • Generalizability: Compared with language models or computer vision, problems from mechanics and engineering design can appear (or sometimes are) overly specific in the eyes of a researcher from, e.g., the machine learning community, even in cases where the domain expert may know of the potentially large impact of a solution of the underlying engineering problem or, more generally, of its generalizability to other problems. Hence, published datasets may remain unused even if they were easily usable and understandable. Nevertheless, researchers in the priority program report from exchanges with data scientists that many seem to express a general interest in challenging and impactful benchmark problems from engineering.
@@ 4 Dealing with the Challenges
Evaluation: If the evaluation mechanism happens to be lightweight and easy to install and use, it can be sufficient to distribute it with the dataset or model problem. In other cases, a workable, yet more involved, solution is to publish a (web-)interface to an evaluator, to which results can be uploaded. The evaluation is performed on the server, and the evaluation results are made available to the user. The advantage is that the server and evaluation toolchain are set up and configured only once by the domain experts. Luckily, the required technology is already available from other domains, albeit for slightly different reasons. In data science, it is sometimes even undesirable to publish test datasets used for evaluation since then researchers may tune their method not only with the training dataset but also with the test dataset. For these reasons, there exists open-source software that greatly helps to implement such online-evaluation or benchmarking tools. A popular example is Codabench [10], which is even used to do public competitions for defined tasks. Some challenges, like the funding and maintenance of the server, remain. However, a growing number of success stories shows that this can work and can significantly help to popularize a benchmark problem, particularly in the crucial phase when the problem is new.
@@
Interpretability and Domain Knowledge: Some of the issues involved are appropriately dealt with by established principles for data management, such as the FAIR principles (short for findable, accessible, interoperable, reusable, [1]), where some of those principles can be fulfilled by including descriptive metadata with the data. There exist professionally curated data repositories implementing these principles, some operated by universities (e.g., the DaRUS repository [13]) or by third parties. There also exist registries of research data repositories, e.g., re3data [14]. For engineering and engineering design problems, mere metadata, such as describing what is stored in which format, may still not be enough to facilitate usage, especially for researchers from other domains. Hence, at the very least, datasets should come shipped with some open-source code that, as a form of minimal example, reads the data from the data files and visualizes it in an intuitive way. The latter is very important in the field of mechanics and dynamics, as a good visualization of the data can be quite insightful – compare, e.g., an explanatory spatial animation of motion to just supplying raw data or line plots in a variety of different coordinate frames. If data and program code are published at different places, hyperlinks should point from one to the other, as well as to and from associated research articles. A further way to increase accessibility and reduce the inevitable technical hurdle to look at published code and data is technology such as Google Colab [15], which permits very straightforward (Python) code execution via a web browser, including access to GPUs. A good explanatory worksheet, which outlines the idea behind the problem, interactively visualizes the data, and explains the basic usage of the proposed methodology, can help greatly to interest and make familiar interested parties with a specialized problem, as the execution of the worksheet is just one hyperlink and one click away. Technology, regulations, and guidelines for research data management advance swiftly. Even binding legislation may appear in the future [16]. Many universities now have strategies for the management of research data that provide useful information. Moreover, helpful resources include the German Research Foundation’s comments on the topic [17], including possibilities to receive funding for data management, which supplement the usual guidelines for good scientific practice [18].
@@ 5 Examples: Published Code and Data
Subsequently, published data and code from within the priority program are listed and discussed, to serve as inspiration, and to learn which types of data and problems appear in mechanics, dynamics, and engineering design.
@@ 5.1 Vibrating Plates
Figure 1: Example plates from the vibrating plates dataset along with the averaged mean squared velocity of the vibrations. Indentations are marked in white in the top row.
@@ 5.2 Database of Crashworthiness Analysis of Crash Boxes
Figure 2: Crash box position in automobiles.
Every year, approximately 2.54 million people in Germany alone are involved in automotive crashes, which makes vehicle crashworthiness a crucial part of the development of automobiles [25]. To help reduce the seriousness of the impact and prevent serious injuries and fatalities, both active and passive crash management systems are installed in automobiles. Passive safety systems are usually structural components which are built to absorb impact energies and reduce potential harm to the occupants. Since these systems are static, they must be meticulously designed and optimized to perform in various scenarios. One such passive safety system is a crash box behind the front bumper, as shown in Figure 2. Crash boxes are thin-walled hollow structures that are installed between the subframe and the front bumper. These absorb the kinetic energy from the impact by undergoing axial deformation and local buckling, thereby reducing the forces experienced by the occupants. The optimal design of the crash box is crucial since frontal impacts are the most common impact type in road traffic.
@@
Figure 3: Pareto front for the objectives: maximum deformed length and total energy absorbed.
Other parameters, such as maximum deformed length, mean contact force, and mass of the crash box, are informative parameters of the crash box’s overall performance. Assuming the energy absorbed has to be maximized while reducing the maximum deformed length following non-dominated solutions, as shown in Figure 3, is obtained. This is called the Pareto front, and it denotes the solutions where both objectives are important.
@@ 5.3 Backbone Reconstruction of a Non-slender Soft Robot
Figure 4: Images of a non-slender soft robot annotated with the reconstructed backbone (blue dashed line) and estimate points (white markers). Taken from [29] by CC BY.
@@ 5.4 Duffing Oscillator Response Analysis (DORA)
Figure 5: DORA time series prediction task in the small-data limit: the training data (left) comprises two trajectories of period-2 cycle dynamics for two values of external forcing amplitudes. The modeling tasks aims at generalization to different forcing amplitudes that induce qualitatively different dynamics, among others chaotic dynamics, as displayed in the right.
@@ 5.5 Passive One-Legged Hopper
In periodic tasks, such as legged locomotion, the resulting motion is often regarded on the basis of a recurrence map, which computes how the state is updated during one full period. As a consequence, the resulting simulations are often computationally expensive. This holds in particular, when gradient information must be computed within the simulation to obtain sensitivities. This derivative information is of great importance not only for the study of the bifurcation behavior, but also for understanding stability and developing efficient and robust controllers for robotic systems. Despite the high complexity involved in generating time-series data for non-smooth dynamical systems, the resulting dataset can still be low dimensional, which makes it an interesting candidate for learning surrogate functions.
@@
Figure 6: Depicted are the passive one-legged hopper and a bifurcation diagram of its periodic trajectories. Since the system conserves energy, the bifurcation parameter is the energy level $\bar{E}$. The projections of the periodic motions presented are the period time $T$ and horizontal velocity $\dot{x}_{0}$ at the apex of the hopper. The illustrations are taken from Figure 3 and 4 in [40].
@@ 5.6 Motion Data of a Four-Bar Mechanism
Figure 7: Hardware prototype of a four-bar mechanism used to stir a viscous fluid, where the stirrer can be replaced. An exemplary stirrer is shown in the lower right.
Figure 8: (a) Prescribed current trajectory and corresponding current measurements. (b) Measured velocity for the applied currents.
By mounting a stirrer onto the mechanism and moving it through a viscous liquid, different damping effects can be introduced into the system. The complete dataset, a notebook for intuitive handling of the data, and a detailed description of the mechanism are published in a curated data repository of the University of Stuttgart [44]. They were used by a partner research group to compare two different hybrid-modeling approaches in [45].

### v2

@@ Abstract
Data-based methods have gained increasing importance in engineering, especially but not only driven by successes with deep artificial neural networks. Success stories are prevalent, e.g., in areas such as data-driven modeling, control and automation, as well as surrogate modeling for accelerated simulation. Beyond engineering, generative and large-language models are increasingly helping with tasks that, previously, were solely associated with creative human processes. Thus, it seems timely to seek artificial-intelligence-support for engineering design tasks to automate, help with, or accelerate purpose-built designs of engineering systems, e.g., in mechanics and dynamics, where design so far requires a lot of specialized knowledge. However, research-wise, compared to established, predominantly first-principles-based methods, the datasets used for training, validation, and test become an almost inherent part of the overall methodology. Thus, data publishing becomes just as important in (data-driven) engineering science as appropriate descriptions of conventional methodology in publications in the past. This article analyzes the value and challenges of data publishing in mechanics and dynamics, in particular regarding engineering design tasks, showing that the latter raise also challenges and considerations not typical in fields where data-driven methods have been booming originally. Possible ways to deal with these challenges are discussed and a set of examples from across different design problems shows how data publishing can be put into practice. The analysis, discussions, and examples are based on the research experience made in a priority program of the German research foundation focusing on research on artificially intelligent design assistants in mechanics and dynamics.
@@ 1 Introduction
Secondly, the article presents a set of seven examples from the aforementioned research projects that outline how the described challenges of data publication were overcome. The authors hope that these concrete examples will further nurture ongoing discussions and change within the community and will be of help to other scientists in mechanics and dynamics who want to publish data and data-driven research.
Thirdly, while not the immediate goal, the article and its examples may show to scientists from outside engineering that very challenging and worthwhile problems exist in mechanics, dynamics, and engineering design, where good data-driven contributions could lead to sustained technological progress. The authors invite data scientists to try their methods on the published data.
@@
At the same time, an increasing number of methodological works exist that deal with generative design in engineering, see, e.g., the review paper [8], giving an overview of various generative methods used for engineering design. Therein, the authors identify several challenges for deep-learning based design methods, including for instance whether deep-learning based approaches can actually be creative enough to produce novel results as well as the usability and feasibility of resulting designs, e.g., for production. Unlike the present work, [8] is focused on the learning methods themselves, not on the aspect of data publishing, its challenges, and implications for the proceedings of science on generative engineering design in mechanics and dynamics. A review of machine-learning methods commonly used in computational mechanics, not necessarily for design tasks, is given by [9]. In general, the breadth of work and methodology discussed in [8, 9] can be seen as a motivation for this work’s approach to look at data-driven design in mechanics and dynamics from a data-centric instead of a method-focused or application-centric perspective. In particular, the article focuses on what that entails for publishing practices.
The article is organized as follows. Based on a general motivation in Section 2, challenges for publishing data in mechanics, dynamics and for associated design tasks are formulated in Section 3. Ideas and guidelines how these challenges might be addressed are discussed in Section 4. Section 5 presents examples of published datasets, program code, and model problems, showing first attempts to meet and overcome the aforementioned challenges, and concludes with lessons learned from the examples. Finally, Section 6 draws conclusions and gives and outlook with regard to the findings specific for engineering design.
@@ 2 The Value of Publishing Data
It is to be expected that the availability of data and well-prepared benchmark problems can very much influence and steer in which areas progresses in machine learning are made. However, an analysis of published datasets and openly accessible and executable benchmark problems shows that there are comparatively few from the field of mechanics and dynamics, especially when focusing on published material suitable for people that are not experts in the field.
Obviously, apart from the kind of data mentioned before, there have always been published datasets that are not meant to be interesting primarily due to their value as a basis for machine learning. For instance, measurements of temperatures at a certain geographic location or measurements of certain material properties are among these datasets for which the provision of the data itself is, at least classically, the purpose of the investigation and publication. This kind of data publishing is not the main purpose of this paper.
@@ 3 Challenges in Data Publishing
Some challenges to be faced for the publication of data are quite common beyond the specific characteristics of the research on design assistant systems. These include:
  • Interpretability and Domain Knowledge: Raw data can be tough to understand for anybody unfamiliar with the generation process of the data. Furthermore, the representation of the data may heavily depend not only on the underlying system but also on the chosen analysis method and observation approach. In simulations, the deformation of a simple mechanical beam, for example, can be computed using numerical discretization schemes that involve the partitioning of the physical domain into small elements. The shape and/or numbering of those elements may however greatly vary across different modeling or simulation approaches, while displaying the same or a very similar elastic deflection of the beam. Comparability of different simulation techniques – and the resulting data – can be difficult for the very same physical system. This can happen analogously in experiments, if measurements are taken at different points of the same mechanical object or with different measurement hardware, which may come with different built-in filtering, e.g., for noise reduction. In the light of this challenge, it is maybe not surprising that many of the first, renewed successes of machine learning in recent years happened for data interpretable without special education, like text, speech, and image data.
@@
  • Generalizability: Compared with language models or computer vision, problems from mechanics and engineering design and corresponding collected data can appear (or sometimes are) overly specific in the eyes of a researcher from, e.g., the machine learning community, even in cases where the domain expert may know of the potentially large impact of a solution of the underlying engineering problem or, more generally, of its generalizability to other problems. Hence, published datasets may remain unused even if they were easily usable and understandable.
@@ 4 Dealing with the Challenges
Evaluation: If the evaluation mechanism happens to be lightweight and easy to install and use, it can be sufficient to distribute it with the dataset or model problem. In other cases, a workable, yet more involved, solution is to publish a (web-)interface to an evaluator, to which results can be uploaded. The evaluation is performed on the server, and the evaluation results are made available to the user. The advantage is that the server and evaluation toolchain are set up and configured only once by the domain experts. Luckily, the required technology is already available from other domains, albeit for slightly different reasons. In data science, it is sometimes even undesirable to publish test datasets used for evaluation since then researchers may tune their method also with the test dataset. For these reasons, there exists open-source software that greatly helps to implement such online-evaluation or benchmarking tools. A popular example is Codabench [12], which is even used to do public competitions for defined tasks. Some challenges, like the funding and maintenance of the server, remain. However, a growing number of success stories shows that this can work and can significantly help to popularize a benchmark problem, particularly in the crucial phase when the problem is new.
@@
Interpretability and Domain Knowledge: Some of the issues involved are appropriately dealt with by established principles for data management, such as the FAIR principles (short for findable, accessible, interoperable, reusable, [1]), where some of those principles can be fulfilled by including descriptive metadata with the data. There exist professionally curated data repositories implementing these principles, some operated by universities (e.g., the DaRUS repository, see [15]) or by third parties. There also exist registries of research data repositories, e.g., re3data [16]. For engineering and engineering design problems, mere metadata, such as describing what is stored in which format, may still not be enough to facilitate usage, especially for researchers from other domains. Hence, at the very least, datasets should come shipped with some open-source code that, as a form of minimal example, reads the data from the data files and visualizes it in an intuitive way. The latter is very important in the field of mechanics and dynamics, as a good visualization of the data can be quite insightful – compare, e.g., an explanatory spatial animation of motion to just supplying raw data or line plots in a variety of different coordinate frames. If data and program code are published at different places, hyperlinks should point from one to the other, as well as to and from associated research articles. A further way to increase accessibility and reduce the inevitable technical hurdle to look at published code and data is technology such as Google Colab [17], which permits very straightforward (Python) code execution via a web browser, including access to GPUs. An explanatory worksheet, which outlines the idea behind the problem, interactively visualizes the data, and explains the basic usage of the proposed methodology, can help greatly to interest and make familiar interested parties with a specialized problem, as the execution of the worksheet is just one hyperlink and one click away. Technology, regulations, and guidelines for research data management advance swiftly. Even binding legislation may appear in the future [18]. Many universities now have strategies for the management of research data that provide useful information. Moreover, helpful resources include the German Research Foundation’s comments on the topic [19], including possibilities to receive funding for data management, which supplement the usual guidelines for good scientific practice [20].
@@ 5 Examples: Published Code and Data
Subsequently, published data and code, as they have resulted from the priority program ”Daring More Intelligence – Design Assistants in Mechanics and Dynamics” (SPP 2353) of the German Research Foundation, are listed and discussed in the light of the aforementioned findings, to serve as inspiration how some of the formulated challenges can be overcome, and to learn which types of data and problems appear in mechanics, dynamics, and engineering design.
@@ 5.1 Vibrating Plates
Figure 1: Example plates from the vibrating plates dataset along with the averaged mean squared velocity of the vibrations. Indentations are marked in white in the top row
@@ 5.2 Database of Crashworthiness Analysis of Crash Boxes
Figure 2: Crash box position in automobiles
Every year, approximately 2.54 million people in Germany alone are involved in automotive crashes, which makes vehicle crashworthiness a crucial part of the development of automobiles [27]. To help reduce the seriousness of the impact and prevent serious injuries and fatalities, both active and passive crash management systems are installed in automobiles. Passive safety systems are usually structural components which are built to absorb impact energies and reduce potential harm to the occupants. Since these systems are static, they must be meticulously designed and optimized to perform in various scenarios. One such passive safety system is a crash box behind the front bumper, as shown in Figure 2. The above model is derived from the Honda Accord Model provided in the NHTSA database [28]. Crash boxes are thin-walled hollow structures that are installed between the subframe and the front bumper. These absorb the kinetic energy from the impact by undergoing axial deformation and local buckling, thereby reducing the forces experienced by the occupants. The optimal design of the crash box is crucial since frontal impacts are the most common impact type in road traffic.
@@
Figure 3: Pareto front for the objectives: maximum deformed length and total energy absorbed
Other parameters, such as maximum deformed length, mean contact force, and mass of the crash box, are informative parameters of the crash box’s overall performance. Assuming the energy absorbed has to be maximized while reducing the maximum deformed length leads to certain non-dominated solutions, as shown in Figure 3. These form the so-called Pareto front, and it denotes the solutions where both objectives are important.
@@ 5.3 Backbone Reconstruction of a Non-slender Soft Robot
Figure 4: Images of a non-slender soft robot annotated with the reconstructed backbone (blue dashed line) and estimate points (white markers). Taken from [32] by CC BY
@@ 5.4 Duffing Oscillator Response Analysis (DORA)
Figure 5: DORA time series prediction task in the small-data limit: the training data (left) comprises two trajectories of period-2 cycle dynamics for two values of external forcing amplitudes. The modeling tasks aims at generalization to different forcing amplitudes that induce qualitatively different dynamics, among others chaotic dynamics, as displayed in the right
@@ 5.5 Passive One-Legged Hopper
In periodic tasks, such as legged locomotion, the resulting motion is often regarded on the basis of a recurrence map, which computes how the state is updated during one full period. As a consequence, the resulting simulations are often computationally expensive. This holds in particular when gradient information must be computed within the simulation to obtain sensitivities. This derivative information is of great importance not only for the study of the bifurcation behavior, but also for understanding stability and developing efficient and robust controllers for robotic systems. Despite the high complexity involved in generating time-series data for non-smooth dynamical systems, the resulting dataset can still be low dimensional, which makes it an interesting candidate for learning surrogate functions.
@@
Figure 6: Depicted are the passive one-legged hopper and a bifurcation diagram of its periodic trajectories. Since the system conserves energy, the bifurcation parameter is the energy level $\bar{E}$. The projections of the periodic motions presented are the period time $T$ and horizontal velocity $\dot{x}_{0}$ at the apex of the hopper. The illustrations are taken from Figure 3 and 4 in [43]
@@ 5.6 Motion Data of a Four-Bar Mechanism
Figure 7: Hardware prototype of a four-bar mechanism used to stir a viscous fluid, where the stirrer can be replaced. An exemplary stirrer is shown in the lower right
Figure 8: (a) Prescribed current trajectory and corresponding current measurements. (b) Measured velocity for the applied currents
By mounting a stirrer onto the mechanism and moving it through a viscous liquid, different damping effects can be introduced into the system. The complete dataset, a notebook for intuitive handling of the data, and a detailed description of the mechanism are published in a curated data repository of the University of Stuttgart [47]. They were used by a partner research group to compare two different hybrid-modeling approaches in [48]. Such an identified model can then be used, e.g., to obtain a model-based controller, furnishing, together with the data-driven modeling process, a control-design assistant.
@@
5.7 Axial Turbine Dataset
In recent years, the demand for renewable energies has drastically grown to reduce the dependency on fossil fuels. Among all, hydropower itself contributes 47% of total renewable energies produced in 2023 [49]. Axial turbines are one of the popular machines for hydropower plants due to their nature to be suitable for low-head conditions. Typically, hydraulic machines are designed for an operating range with several operating points simultaneously. In their simplest form, they are designed specifically for one operating condition to be fulfilled. For hydraulic machines, the common objectives are design head ($\Delta h$; energy difference between flow inlet and outlet of the turbine), efficiency ($\eta$), power ($P$), mechanical stresses ($\sigma$) and volume of cavitation ($V_{\textnormal{Cav}}$; the volume of fluid (vapor-filled cavities) below the vapor pressure within the liquid). Hence, multi-objective optimizations are performed to get a finalized shape of the axial turbines [50]. Performing multi-objective optimization of an axial turbine is expensive due to the number of expensive computational fluid dynamics (CFD) simulations involved.
(a) Axial turbine test case
(b) Blade parameterization
Figure 9: Section of an axial turbine and the parameterization of the blade [51, 52]
Figure 9 visualizes a typical section of an axial machine and the corresponding b-splines to parameterize the blade. The detailed descriptions of the axial turbine and the parameterization are published in [51, 52]. Usually, hydraulic turbines are operated at different operating conditions like part, full, and nominal-loads. Hence during optimization, it is important to take care of operating conditions as well. This means that for each operating point, a separate CFD simulation has to be performed. The overarching optimization target is to obtain high-efficiency values and no cavitating regions for the desired net head and power range.
The dataset presented here is an outcome of an optimization run of an axial turbine shown in Figure 9. The inputs of the dataset consist of 30 geometry parameters, which define the blade at the hub, shroud, and mid-span. The shroud and hub diameter remain constant in the current case. The outputs of the dataset are the corresponding post-processed results of the following CFD simulations. They are power ($P$), design head ($\Delta h$), the efficiency of the machine ($\eta$) and volume of cavitation ($V_{\textnormal{Cav}}$). The presented dataset is publicly accessible at the Github repository [53] containing geometry and result data in tabular form. Proper descriptions of inputs are given in the open-access article [52]. It is also possible to generate and extend the dataset. The expansion of the dataset for training, validation, and testing will include CFD simulations to perform, which is computationally expensive. The docker container is available with all source code on another Github repository, which is based on the design tool (d)esign (t)ool (O)bject-(O)riented (dtOO) [54, 55]. Figure 10 illustrates the trade-off between different objective functions. The presented Pareto front in Figure 10(a) is a two-dimensional projection of the three-dimensional set of non-dominated solutions, displaying the key objectives when the volume of cavitation is minimized at the same time, maximizing the efficiency of the machine.
This comprehensive axial turbine dataset offers valuable resources for both engineers and machine learning practitioners. Engineers can utilize this data to optimize turbine designs, and benchmark performance across various operational conditions. The inclusion of cavitation data is particularly useful for improving turbine longevity and reliability. Machine learning engineers can leverage this dataset to develop predictive models for turbine performance, potentially creating algorithms that can forecast efficiency and cavitation risks based on input parameters. Such models could be invaluable for real-time monitoring and predictive maintenance in hydropower plants. Additionally, the dataset could be used to train deep learning models for automated design optimization, potentially uncovering novel turbine configurations that maximize efficiency while minimizing cavitation.
(a) The results present the volume of cavitation versus the efficiency of the machine at nominal load
(b) The results present the volume of cavitation versus the efficiency versus design head of the machine at nominal load.
Figure 10: Pareto front of the dataset after 89 generations
The dataset of this relatively simple turbine geometry is interpretable with basic fluid mechanics and turbomachinery knowledge. It is available to train machine-learning-based algorithms, specifically mapping from different design parameters to the post-processed results. Regarding generalizability, since hydraulic machines are specifically designed for specific applications, the design or machine type cannot be generalized in the sense of large-scale production. However, this simple case may serve as a training object to optimize more complex turbine types in further work. Furthermore, the prediction of the flow field is based on the numerical simulation of the coupled nonlinear, partial differential equations (Navier-Stokes equations), which makes the generalizability in the case of hydraulic turbines impossible. In terms of the problem complexity, the multi-objective problem can be changed to a single-objective optimization. Neglecting other objectives can significantly reduce the complexity of the dataset. However, with the current state-of-the-art turbine design problems, the complexity will substantially increase, primarily because of an increase in the number of geometry parameters. Owing to the flow interaction between several turbine components, the optimization of just one component is not sufficient. On the contrary, it is essential to design several turbine components simultaneously in order to optimize the entire machine.
5.8 Lessons Learned
While translating the perceived challenges and proposed solution approaches of data publishing in the studied area to the seven concrete examples, the authors learned various lessons that can be summarized as follows, focusing on aspects not already discussed at length:
  • Researchers seem to routinely underestimate the time and effort it takes for others to use their data and code.
  • A review process for data publications and associated code is usually necessary to ensure quick accessibility even for other domain experts. The reviewer needs to actually execute the software and load and work with the data to ensure its functionality on an unprepared computer. As this is not usually required for publications of any sort, one should ask fellow scientists to try it and report back how long it took them to make it work. This practice was used here where applicable.
  • It can indeed be non-trivial to manually or automatically evaluate the quality, e.g., of design results in mechanics and dynamics, different from, e.g., applications of learning methods to natural language and images.
  • All authors have worked with their own published datasets described here and partly with those of their co-authors to use them within engineering design. Many have experienced that the composition of the data and potential data preprocessing or alternative data representations were usually very fundamental to the performance of the developed approach. For many engineering tasks, it is still very unclear which data representation is best suited, unlike for some other fields like, e.g., computer vision where successful ways to represent images for learning have been established. Sometimes, it is even unclear which data of an engineering system to record, to begin with.
  • Pure publications of datasets in a data repository are nowadays citable with a digital object identifier. However, for key performance indicators used to assess researchers’ careers, such as the $h$-index, data publications of this kind and citations of them are often irrelevant. This is in contrast to the undisputed importance of data for data-based methods and the skill and effort it requires to obtain and prepare a useful dataset. This discourages the publication of data for which one is not publishing a paper in parallel that uses this data and neglects the worth of published data in its own right. As papers usually propose functioning solutions, in effect, publication of data for unsolved problems is discouraged.